OpenStudy (anonymous):

use the substitution method to solve the given system of equations. express non-integer values as fractions in lowest terms. 4x + 2y=7 y=5x a.(2/5,2) b.(7/11,35/11) c.(1/2,5/2) d.(2,10)

OpenStudy (anonymous):

C

OpenStudy (campbell_st):

substitute equation 2 into equation 1 4x + 2(5x) = 7 14x = 7 x = 7/14 or 1/2 substitute this value into equation 2 y = 5(1/2) y = 5/2 the ordered pair is (1/2, 5/2)